6 of 6 (6 complete) HW Score: 83.33%, 5 of 6 pts Score: 0 of 1 pt X 10.3.17-T Question Help Listed below are altitudes (thous a. find the explained variation
b.find the unexplained variation
c.find the indicated prediction interval
Answer #1

Consider altitude be the independent variable, x and temperature be the dependent variable, y. Minitab is used to analyze theRegression Equation Temperature = 68.98 - 3.485 Altitude Coefficients Coef SE Coef T-Value P-Value Term 95% CI VIF 68.98 6.72(b) From the obtained output, the unexplained variation is, SSE = E(y-i) = 355.7 (c) The 95% prediction interval when altitudRegression Equation Temperature 68.98 - 3.485 Altitude !! Settings Variable Setting Altitude 6.327 Prediction Fit SE Fit 95%